# Better Call SAL: Towards Learning to Segment Anything in Lidar

### v1

@@ A.3 Text Prompt Engineering
Table A.2: CLIP token distillation and text prompt engineering. To evaluate our token prediction, we prompt the SemanticKITTI class vocabulary to generate labeled training data and train a non-zero-shot model (row 1). Furthermore, we demonstrate the insufficiency of vanilla class names (car) as text prompts and the boost from engineering a rich set of terms (car, jeep, SUV, van) as explained in Sec.˜A.3.
@@
$\times$ $19.6$ $26.0$ $65.8$ $22.1$
$\times$ $\times$ $23.6$ $31.2$ $67.0$ $29.0$
@@
$\times$ $23.2$ $28.6$ $51.5$ $26.1$
$\times$ $\times$ $46.3$ $60.1$ $76.1$ $52.3$
A.4 How to Train on Partial Labels?
Pseudo-labels provide only partial supervision within the camera frustum (see Fig.˜2(a)), leaving the majority of Lidar points unlabeled (see Tab.˜B.1). How can we train SAL with such partial supervision?
Ignore unlabeled region + standard data augmentations.  During training, we remove all unlabeled points from the point cloud. Otherwise, the $\mathcal{L}_{obj}$ loss would penalize any segmentation in these regions as a false positive, thereby teaching the model to ignore them entirely. As visualized in Fig.˜2(a) and quantified in Tab.˜B.1, the 360 Lidar label coverage is particularly low for single camera setups, as in SemanticKITTI [6].
@@ A.4 How to Train on Partial Labels?
Figure A.2: Training on partial labels. Unprojecting image-based pseudo labels results in a partially (pseudo) labeled point cloud (Fig.˜2(a)). We construct supervisory signal by concatenating multiple partially labeled point clouds (Fig.˜2(b)).
Appendix B Pseudo-label Analysis
B.1 Pseudo-label Statistics
Point coverage.  As shown in Tab.˜B.1, the single-camera setup of Semantic-KITTI [6] allows us to label a significantly smaller portion of Lidar point clouds ($14\%$) compared to GT labels ($98\%$). Interestingly, even for nuScenes dataset [18], which provides a setup with five cameras and $360^{\circ}$ view coverage, pseudo-labels cover only $48\%$ of all points. Furthermore, we quantify pseudo-label coverage on SemanticKITTI when filtering all points outside of the camera frustum. This leads to coverage of $89\%$. The remaining missing labels can be explained by errors committed by our image-based segmentation foundation model, SAM [31]. This analysis confirms that even when utilizing strong foundation models, image-to-Lidar distillation remains a challenging problem.
@@ B.1 Pseudo-label Statistics
Table B.1: Pseudo-label statistics. We outline the label coverage of point clouds, the total, max, and mean number of instances per scan, and the ratio of things/stuff instances on the full point cloud and point cloud areas that overlap with the camera view frustum (Filter Frustum). As can be seen, due to the single-camera setup, pseudo-label coverage in SemanticKITTI [6] is very low ($14\%$ of points). Even though nuScenes [18] dataset provides $360^{\circ}$ view coverage, only $48\%$ are labeled due to blind spots. Even when only retaining points, that overlap with the camera view frustum (SemanticKITTI, Filter Frustum), we observe coverage of $89\%$. This can be explained by mistakes (e.g., false negatives) committed by our segmentation foundation model (SAM [31]).
@@
Fine-grained analysis.  In Tab.˜B.2, we additionally report per-class label statistics for train and validation splits for SemanticKITTI dataset [6]. As can be seen, the label coverage is consistent in train and val splits ($97\%$ GT vs. $14\%$ pseudo-labels in val, and $98\%$ GT vs. $15\%$ pseudo-labels in train). Overall, the max. number of instances is larger in the train set ($65$ GT & $122$ pseudo) as compared to the validation set ($53$ GT & $82$ pseudo), whereas the average number of instances remains consistent for pseudo-labels, while for GT labels are lower in the train ($19$ GT & $40$ pseudo) compared to val ($24$ GT & $41$ pseudo). The larger number of instances reflect the most frequent classes: on the train set (Filter Frustum), the highest percentage of instances are due to vegetation ($27.1\%$), building ($16.9\%$), and road ($13.4\%$) classes. As can be confirmed in Fig.˜E.3 and Fig.˜E.4, our pseudo-labels indeed often localize individual trees, bushes, and buildings, leading to a large number of overall instances for these classes.
Table B.2: Pseudo-label statistics per class on SemanticKITTI. We compare ground truth label and pseudo-label statistics on SemanticKITTI [6] train and validation sets. Label coverage and instance class distributions are reported in percentage. Due to camera visibility coverage, pseudo-labels cover a significantly smaller portion of the dataset than GT labels. This is especially prominent in the SemanticKITTI dataset ($14\%$ coverage). Pseudo-labels cover a significantly larger number of instances per scan than GT labels ($3\times$ more when only compared in the camera view frustum), as needed to learn to segment a large variety of objects. While GT labels treat stuff classes as a “single instance,” our pseudo-labels hypothesize a variety of plausible segmentations of stuff classes, leading to a higher percentage of stuff class labels. We use a semantic oracle to report per-class statistics on pseudo-labels
@@
Table B.3: Segmentation pseudo-labels analysis. This table extends Tab.˜4 from the main paper with additional metrics and baselines. We report results on the SemanticKITTI validation set using the semantic oracle, as detailed in the main paper (Sec.˜4.1). We evaluate the region of the point cloud visible in the camera (Filter Frustum).
@@
Table B.4: Segmentation pseudo-labels ablation - per class. This table extends Tab.˜4 from the main paper with per-class PQ scores.. We report results on the SemanticKITTI validation set using the semantic oracle, as detailed in the main paper (Sec.˜4.1). We evaluate the region of the point cloud visible in the camera (Filter Frustum).
@@
B.2 Pseudo-label Timing & Effort Analysis
We report information and statistics on three datasets (SemanticKITTI [6], Panoptic nuScenes [18] and Waymo Open [75]) that we pseudo-label and use to train SAL, in Tab.˜B.5. As can be seen, the processing time needed to process a single scan depends on multiple dataset characteristics, such as the number of cameras. In the case of the Waymo Open dataset, we pseudo-label every $10^{th}$ scan, and we do not perform DBSCAN refinement. The reason is two-fold: (i) we observe image-lidar calibration on Waymo is more accurate compared to KITTI and nuScenes, and (ii) we save processing effort/time to pseudo-label Waymo. Importantly, our pseudo-labeling setup is general: it supports various sensory setups (single camera, multi-camera) and multiple Lidar types and can cope well with datasets with different degrees of accuracy of the image-lidar calibration/synchronization.
In Tab.˜B.6, we provide fine-grained per-scan timing analysis for SAL pseudo-labeling engine. The most costly component is extracting image-level segmentation masks using the segmentation foundation model (SAM [31]), which needs to be performed, in general, for each camera.
Table B.5: Dataset pseudo-labeling analysis. We report dataset information and statistics, along with pseudo-labeling effort and time analysis. We note that in SemanticKITTI, we have four cameras in total. However, only one is used for pseudo-labeling, as all cameras have (roughly) the same field of view. We report the total number of scans, as well as per-scan processing time (NVIDIA A100D-80C GPU). Processing time depends on the number of cameras, as well as individual dataset characteristics (such as camera resolution, point cloud size, density, etc.). Finally, we report the total time needed to pseudo-label a dataset if processed sequentially (in practice, we can pseudo-label datasets in 1-3 days using a compute cluster). We note that in the case of Waymo, due to sheer size, we (i) pseudo-label only $10\%$ of point clouds and (ii) skip the postprocessing with DBSCAN.
@@ B.2 Pseudo-label Timing & Effort Analysis
SemanticKITTI 11 90 Velodyne HDL-64E 43,592/43,59243,\!592/43,\!592 $32$ $16$
Panoptic nuScenes 66 $360$ Velodyne HDL-32E 40,157/40,15740,\!157/40,\!157 $129$ $59$
Waymo Open 55 $270$ Waymo Proprietary 227,101/22,710 92 24
Table B.6: Per-scan pseudo-labeling timing analysis. We report fine-grained per-dataset timing analysis for our pseudo-labeling system. In particular, we report individual timings for core components of SAL pseudo-label engine: (i) running segmentation foundation model (SAM [31]) in the image domain, (ii) extracting corresponding CLIP features [16, 67], (iii) image-to-Lidar unprojection, and finally, (iv) DBSCAN refinement. As can be seen, the most costly step is due to SAM; this step could be reduced in the future using newer and more time-efficient variants of SAM. DBSCAN is the second bottleneck, however, it is only executed once per point cloud, and does not depend on the number of cameras. The analysis was conducted using NVIDIA A100D-80C GPU.
@@ B.3 Qualitative Analysis
Both SemanticKITTI [6] and nuScenes [18] do not provide ground truth annotations in 2D. To illustrate the performance of the foundation models we utilize for pseudo-labeling in the image domain, we visualize predicted masks and dataset class vocabularies from SAM [31] and CLIP [67]. In Fig.˜E.3, we show the single (front) camera view corresponding to the first scan of four different SemanticKITTI sequences. For nuScenes Fig.˜E.4 and Waymo Open, we focus on the first scan of a single sequence and show all camera views. As can be seen, in all datasets, the output contains many correct segmentations and classifications, e.g., for the road, car, or vegetation classes. Moreover, the class-agnostic SAM masks introduce many smaller instances by segmenting individual road markings. However, without any corresponding points in the cloud, such segments are filtered by the unprojection/transfer to Lidar.
@@
While the SAM masks are generally correct, noisy CLIP predictions confirm there is room for improvement. For example, the sky is always misclassified since it is not part of the class vocabulary. As shown in the main paper (Tab.˜3), the distillation of SAM and CLIP to 3D yields an analogous behavior of our SAL model. It should be noted that the clip predictions shown in Fig.˜E.3 and Fig.˜E.4 are only for visualization purposes. We do not directly transfer class labels to 3D, but only the CLIP embeddings.
B.4 Pseudo-label Quality
@@
We additionally report per-class pseudo-label results in Tab.˜B.4. As can be seen, DBSCAN replace consistently improves over tuned SAM variant on things classes (e.g., for car class, +1.6+1.6 PQ, person class +4.8+4.8 PQ). While DBSCAN filter leads to remarkable improvements on some classes (e.g., for car +10.7+10.7, bicycle +43.2+43.2), it severely degrades stuff classes (e.g., 39.8-39.8 for terrain class). On the other hand, DBSCAN replace either does not affect stuff classes or leads to improvements (e.g., building +3.9+3.9).
Appendix C Model Ablations
C.1 Zero-shot Classification
@@
The SAL model performs zero-shot classification by matching text prompts with predicted CLIP tokens. To predict tokens only with Lidar input, our label engine generates pairs of Lidar segments and corresponding CLIP image tokens. Training on these pairs distills the CLIP image encoder into our model. The first row of Tab.˜A.2 demonstrates a different approach where we train on pairs of Lidar segments and class labels obtained by directly prompting the CLIP image encoder. Examples of these labels are visualized in Fig.˜E.3. The difference (1.5-1.5) between the first and third row in Tab.˜A.2 indicates the performance drop of the distillation into our model. This further demonstrates the effectiveness of our framework. In particular, our insufficiencies with respect to zero-shot classification are not caused by our design choices in 3D but merely a reflection of the same limitations in the image domain.
@@
Text prompt engineering.  The CLIP tokens predicted by our model can be prompted with any arbitrary text. This allows our model to classify potentially any object. The text prompts must adhere to the respective class vocabulary to perform zero-shot classification on an annotated dataset. In Tab.˜A.2, we demonstrate the performance boost from engineering the set of text prompts, e.g., by adding similar class terms to each class set (see Tab.˜E.1). In particular, for super-classes, where we add all default class names to the respective super set, our engineering yields a huge gain of +23.1+23.1. This is due to the ambiguous nature of the super class names, e.g., object or structure.
@@
Table C.1: SAL per class PQ results on the SemanticKITTI and nuScenes validation sets. We report metrics for zero-shot (ZS) and linear probing (LP) supervision.
@@
Appendix D Per-class Results
@@
In Tab.˜C.1, we report per-class PQ results of the SAL model on SemanticKITTI [6] and nuScenes [18]. These results correspond to the default class evaluations shown in the main paper (Tab.˜5). For other-X classes which are defined in delimitation to other classes, e.g., other-vehicle than car, our zero-shot model suffers from the ambiguous class prompt (see Sec.˜A.3). Linear probing improves across most classes except, for example, bicycle. The performance drop can be explained by the different object-notion between our pseudo and the ground truth labels.
Appendix E Qualitative Results

### v2

@@ 0.A.3 Text Prompt Engineering
Table 0.A.2: CLIP token distillation and text prompt engineering. To evaluate our token prediction, we prompt the SemanticKITTI class vocabulary to generate labeled training data and train a non-zero-shot model (row 1). Furthermore, we demonstrate the insufficiency of vanilla class names (car) as text prompts and the boost from engineering a rich set of terms (car, jeep, SUV, van) as explained in Sec.˜0.A.3.
@@
$\times$ $20.6$ $27.1$ $65.2$ $20.9$
$\times$ $\times$ $24.8$ $32.3$ $66.8$ $29.7$
@@
$\times$ $27.4$ $33.8$ $71.8$ $24.9$
$\times$ $\times$ $48.5$ $59.4$ $78.8$ $52.8$
0.A.4 How to Train on Partial Labels?
Pseudo-labels provide only partial supervision within the camera frustum (see Fig.˜2(a)), leaving the majority of Lidar points unlabeled (see Tab.˜0.B.1). How can we train SAL with such partial supervision?
Ignore unlabeled region + standard data augmentations.  During training, we remove all unlabeled points from the point cloud. Otherwise, the $\mathcal{L}_{obj}$ loss would penalize any segmentation in these regions as a false positive, thereby teaching the model to ignore them entirely. As visualized in Fig.˜2(a) and quantified in Tab.˜0.B.1, the 360 Lidar label coverage is particularly low for single camera setups, as in SemanticKITTI [6].
@@ 0.A.4 How to Train on Partial Labels?
Figure 0.A.2: Training on partial labels. Unprojecting image-based pseudo labels results in a partially (pseudo) labeled point cloud (Fig.˜2(a)). We construct supervisory signal by concatenating multiple partially labeled point clouds (Fig.˜2(b)).
Appendix 0.B Pseudo-label Analysis
0.B.1 Pseudo-label Statistics
Point coverage.  As shown in Tab.˜0.B.1, the single-camera setup of Semantic-KITTI [6] allows us to label a significantly smaller portion of Lidar point clouds ($14\%$) compared to GT labels ($98\%$). Interestingly, even for nuScenes dataset [18], which provides a setup with five cameras and $360^{\circ}$ view coverage, pseudo-labels cover only $48\%$ of all points. Furthermore, we quantify pseudo-label coverage on SemanticKITTI when filtering all points outside of the camera frustum. This leads to coverage of $89\%$. The remaining missing labels can be explained by errors committed by our image-based segmentation foundation model, SAM [31]. This analysis confirms that even when utilizing strong foundation models, image-to-Lidar distillation remains a challenging problem.
@@ 0.B.1 Pseudo-label Statistics
Table 0.B.1: Pseudo-label statistics. We outline the label coverage of point clouds, the total, max, and mean number of instances per scan, and the ratio of things/stuff instances on the full point cloud and point cloud areas that overlap with the camera view frustum (Filter Frustum). As can be seen, due to the single-camera setup, pseudo-label coverage in SemanticKITTI [6] is very low ($14\%$ of points). Even though nuScenes [18] dataset provides $360^{\circ}$ view coverage, only $48\%$ are labeled due to blind spots. Even when only retaining points, that overlap with the camera view frustum (SemanticKITTI, Filter Frustum), we observe coverage of $89\%$. This can be explained by mistakes (e.g., false negatives) committed by our segmentation foundation model (SAM [31]).
@@
Fine-grained analysis.  In Tab.˜0.B.2, we additionally report per-class label statistics for train and validation splits for SemanticKITTI dataset [6]. As can be seen, the label coverage is consistent in train and val splits ($97\%$ GT vs. $14\%$ pseudo-labels in val, and $98\%$ GT vs. $15\%$ pseudo-labels in train). Overall, the max. number of instances is larger in the train set ($65$ GT & $122$ pseudo) as compared to the validation set ($53$ GT & $82$ pseudo), whereas the average number of instances remains consistent for pseudo-labels, while for GT labels are lower in the train ($19$ GT & $40$ pseudo) compared to val ($24$ GT & $41$ pseudo). The larger number of instances reflect the most frequent classes: on the train set (Filter Frustum), the highest percentage of instances are due to vegetation ($27.1\%$), building ($16.9\%$), and road ($13.4\%$) classes. As can be confirmed in Fig.˜0.E.3 and Fig.˜0.E.4, our pseudo-labels indeed often localize individual trees, bushes, and buildings, leading to a large number of overall instances for these classes.
Table 0.B.2: Pseudo-label statistics per class on SemanticKITTI. We compare ground truth label and pseudo-label statistics on SemanticKITTI [6] train and validation sets. Label coverage and instance class distributions are reported in percentage. Due to camera visibility coverage, pseudo-labels cover a significantly smaller portion of the dataset than GT labels. This is especially prominent in the SemanticKITTI dataset ($14\%$ coverage). Pseudo-labels cover a significantly larger number of instances per scan than GT labels ($3\times$ more when only compared in the camera view frustum), as needed to learn to segment a large variety of objects. While GT labels treat stuff classes as a “single instance,” our pseudo-labels hypothesize a variety of plausible segmentations of stuff classes, leading to a higher percentage of stuff class labels. We use a semantic oracle to report per-class statistics on pseudo-labels
@@
Table 0.B.3: Segmentation pseudo-labels analysis. This table extends Tab.˜4 from the main paper with additional metrics and baselines. We report results on the SemanticKITTI validation set using the semantic oracle, as detailed in the main paper (Sec.˜4.1). We evaluate the region of the point cloud visible in the camera (Filter Frustum).
@@
Table 0.B.4: Segmentation pseudo-labels ablation - per class. This table extends Tab.˜4 from the main paper with per-class PQ scores. We report results on the SemanticKITTI validation set using the semantic oracle, as detailed in the main paper (Sec.˜4.1). We evaluate the region of the point cloud visible in the camera (Filter Frustum).
@@
0.B.2 Pseudo-label Timing & Effort Analysis
We report information and statistics on three datasets (SemanticKITTI [6], Panoptic nuScenes [18] and Waymo Open [75]) that we pseudo-label and use to train SAL, in Tab.˜0.B.5. As can be seen, the processing time needed to process a single scan depends on multiple dataset characteristics, such as the number of cameras. In the case of the Waymo Open dataset, we pseudo-label every $10^{th}$ scan, and we do not perform DBSCAN refinement. The reason is two-fold: (i) we observe image-lidar calibration on Waymo is more accurate compared to KITTI and nuScenes, and (ii) we save processing effort/time to pseudo-label Waymo. Importantly, our pseudo-labeling setup is general: it supports various sensory setups (single camera, multi-camera) and multiple Lidar types and can cope well with datasets with different degrees of accuracy of the image-lidar calibration/synchronization.
In Tab.˜0.B.6, we provide fine-grained per-scan timing analysis for SAL pseudo-labeling engine. The most costly component is extracting image-level segmentation masks using the segmentation foundation model (SAM [31]), which needs to be performed, in general, for each camera.
Table 0.B.5: Dataset pseudo-labeling analysis. We report dataset information and statistics, along with pseudo-labeling effort and time analysis. We note that in SemanticKITTI, we have four cameras in total. However, only one is used for pseudo-labeling, as all cameras have (roughly) the same field of view. We report the total number of scans, as well as per-scan processing time (NVIDIA A100D-80C GPU). Processing time depends on the number of cameras, as well as individual dataset characteristics (such as camera resolution, point cloud size, density, etc.). Finally, we report the total time needed to pseudo-label a dataset if processed sequentially (in practice, we can pseudo-label datasets in 1-3 days using a compute cluster). We note that in the case of Waymo, due to sheer size, we (i) pseudo-label only $10\%$ of point clouds and (ii) skip the postprocessing with DBSCAN.
@@ 0.B.2 Pseudo-label Timing & Effort Analysis
SemanticKITTI 11 90 Velodyne HDL-64E $43592/43592$ $32$ $16$
Panoptic nuScenes 66 $360$ Velodyne HDL-32E $40157/40157$ $129$ $59$
Waymo Open 55 $270$ Waymo Proprietary $227101/22710$ $92$ $24$
Table 0.B.6: Per-scan pseudo-labeling timing analysis. We report fine-grained per-dataset timing analysis for our pseudo-labeling system. In particular, we report individual timings for core components of SAL pseudo-label engine: (i) running segmentation foundation model (SAM [31]) in the image domain, (ii) extracting corresponding CLIP features [16, 67], (iii) image-to-Lidar unprojection, and finally, (iv) DBSCAN refinement. As can be seen, the most costly step is due to SAM; this step could be reduced in the future using newer and more time-efficient variants of SAM. DBSCAN is the second bottleneck, however, it is only executed once per point cloud, and does not depend on the number of cameras. The analysis was conducted using NVIDIA A100D-80C GPU.
@@ 0.B.3 Qualitative Analysis
Both SemanticKITTI [6] and nuScenes [18] do not provide ground truth annotations in 2D. To illustrate the performance of the foundation models we utilize for pseudo-labeling in the image domain, we visualize predicted masks and dataset class vocabularies from SAM [31] and CLIP [67]. In Fig.˜0.E.3, we show the single (front) camera view corresponding to the first scan of four different SemanticKITTI sequences. For nuScenes Fig.˜0.E.4 and Waymo Open, we focus on the first scan of a single sequence and show all camera views. As can be seen, in all datasets, the output contains many correct segmentations and classifications, e.g., for the road, car, or vegetation classes. Moreover, the class-agnostic SAM masks introduce many smaller instances by segmenting individual road markings. However, without any corresponding points in the cloud, such segments are filtered by the unprojection/transfer to Lidar.
@@
While the SAM masks are generally correct, noisy CLIP predictions confirm there is room for improvement. For example, the sky is always misclassified since it is not part of the class vocabulary. As shown in the main paper (Tab.˜3), the distillation of SAM and CLIP to 3D yields an analogous behavior of our SAL model. It should be noted that the clip predictions shown in Fig.˜0.E.3 and Fig.˜0.E.4 are only for visualization purposes. We do not directly transfer class labels to 3D, but only the CLIP embeddings.
0.B.4 Pseudo-label Quality
@@
We additionally report per-class pseudo-label results in Tab.˜0.B.4. As can be seen, DBSCAN replace consistently improves over tuned SAM variant on things classes (e.g., for car class, +1.6+1.6 PQ, person class +4.8+4.8 PQ). While DBSCAN filter leads to remarkable improvements on some classes (e.g., for car +10.7+10.7, bicycle +43.2+43.2), it severely degrades stuff classes (e.g., 39.8-39.8 for terrain class). On the other hand, DBSCAN replace either does not affect stuff classes or leads to improvements (e.g., building +3.9+3.9).
Appendix 0.C Model Ablations
0.C.1 Zero-shot Classification
@@
The SAL model performs zero-shot classification by matching text prompts with predicted CLIP tokens. To predict tokens only with Lidar input, our label engine generates pairs of Lidar segments and corresponding CLIP image tokens. Training on these pairs distills the CLIP image encoder into our model. The first row of Tab.˜0.A.2 demonstrates a different approach where we train on pairs of Lidar segments and class labels obtained by directly prompting the CLIP image encoder. Examples of these labels are visualized in Fig.˜0.E.3. The difference (1.5-1.5) between the first and third row in Tab.˜0.A.2 indicates the performance drop of the distillation into our model. This further demonstrates the effectiveness of our framework. In particular, our insufficiencies with respect to zero-shot classification are not caused by our design choices in 3D but merely a reflection of the same limitations in the image domain.
@@
Text prompt engineering.  The CLIP tokens predicted by our model can be prompted with any arbitrary text. This allows our model to classify potentially any object. The text prompts must adhere to the respective class vocabulary to perform zero-shot classification on an annotated dataset. In Tab.˜0.A.2, we demonstrate the performance boost from engineering the set of text prompts, e.g., by adding similar class terms to each class set (see Tab.˜0.E.1). In particular, for super-classes, where we add all default class names to the respective super set, our engineering yields a huge gain of +23.1+23.1. This is due to the ambiguous nature of the super class names, e.g., object or structure.
@@
Table 0.C.1: SAL per class PQ results on the SemanticKITTI and nuScenes validation sets. We report metrics for zero-shot (ZS) and linear probing (LP) supervision.
@@
Appendix 0.D Per-class Results
@@
In Tab.˜0.C.1, we report per-class PQ results of the SAL model on SemanticKITTI [6] and nuScenes [18]. These results correspond to the default class evaluations shown in the main paper (Tab.˜5). For other-X classes which are defined in delimitation to other classes, e.g., other-vehicle than car, our zero-shot model suffers from the ambiguous class prompt (see Sec.˜0.A.3). Linear probing improves across most classes except, for example, bicycle. The performance drop can be explained by the different object-notion between our pseudo and the ground truth labels.
Appendix 0.E Qualitative Results